# Formatting Instructions For NeurIPS 2024

### v1

@@ Abstract
The recently proposed Bayesian Flow Networks (BFNs) show great potential in modeling parameter spaces, offering a unified strategy for handling continuous, discretized, and discrete data. However, BFNs cannot learn high-level semantic representation from the parameter space since common encoders, which encode data into one static representation, cannot capture semantic changes in parameters. This motivates a new direction: learning semantic representations hidden in the parameter spaces to characterize mixed-typed noisy data. Accordingly, we propose a representation learning framework named ParamReL, which operates in the parameter space to obtain parameter-wise latent semantics that exhibit progressive structures. Specifically, ParamReL proposes a self-encoder to learn latent semantics directly from parameters, rather than from observations. The encoder is then integrated into BFNs, enabling representation learning with various formats of observations. Mutual information terms further promote the disentanglement of latent semantics and capture meaningful semantics simultaneously. We illustrate conditional generation and reconstruction in ParamReL via expanding BFNs, and extensive quantitative experimental results demonstrate the superior effectiveness of ParamReL in learning parameter representation.
@@ 1 Introduction
Representation learning [4], which aims at learning low-dimensional latent semantics from high-dimensional observations, offers an unsupervised approach to discovering high-level semantics in observations. It has been widely applied in areas such as computer vision [26, 53, 10], and data analytics [41, 30]. While most representation learning methods [21, 6, 29] work on continuous-valued observations, different non-trivial methods are needed to discover semantics for the discretized  [44, 32] and discrete data [2, 7]. Consequently, these individual efforts might face issues such as inconsistent discoveries within the data [57] or repeated modelling efforts [23, 54].
On the other hand, Bayesian Flow Networks (BFNs) [16, 38, 50] have been recently proposed as a promising deep generative model. By operating in the parameter space, BFNs design a multi-step mechanism to approximate the ground-truth parameters of observation sequentially. As a result, a uniform strategy may be adopted to deal with continuous, discretized, and discrete data while simultaneously maintaining fast sampling. Pilot studies of BFNs have shown promising results in modelling different data formats.
@@ 2 Bayesian Flow Networks
Bayesian Flow Networks (BFNs) [16, 38, 50] are a new class of latent variable models. Similar to the reverse process of diffusion models [18, 37], BFNs target optimizing the neural network architecture on posterior parameters to simulate observations as they are gradually observed. In particular, BFNs assume two types of distributions: a simple input distribution $P_{\mathrm{I}}(\cdot)$ representing the initial belief about observations and an output distribution $P_{\mathrm{O}}(\cdot)$ representing the simulating distribution for observations. The parameters of input distribution are first updated through a Bayesian inference scheme and then passed into a neural network to form the parameters of output distributions. The main objective of BFNs is to minimize the Kullback-Leibler (KL) divergence between the ground-truth data distribution and the output distribution, ensuring that the output distribution closely approximates the ground-truth data distribution.
@@
Following the notations in diffusion models, we use ${\boldsymbol{\mathbf{x}}}_{0}$ to denote the observations. There are $T$ reverse steps in BFNs which gradually reveals the information of ${\boldsymbol{\mathbf{x}}}_{0}$ through $\{{\boldsymbol{\mathbf{x}}}_{T},{\boldsymbol{\mathbf{x}}}_{T-1},\ldots,{\boldsymbol{\mathbf{x}}}_{1}\}$ to the input distribution [^1] It is noted that the index $t$ is used reversely in [16]. We make such changes to be consistent with the diffusion model settings [18, 37].. At the $t$-th step, the reverse step is denoted as a sender distribution $p_{\mathrm{S}}({\boldsymbol{\mathbf{x}}}_{t-1}\,|\,{\boldsymbol{\mathbf{x}}}_{t};\alpha_{t})$, where $\alpha_{t}$ is a precision parameter. Combined with input distribution $p_{\mathrm{I}}({\boldsymbol{\mathbf{x}}}_{t};{\boldsymbol{\mathbf{\theta}}}_{t+1})$, the posterior distribution of ${\boldsymbol{\mathbf{x}}}_{t}$ is obtained as $p({\boldsymbol{\mathbf{x}}}_{t};{\boldsymbol{\mathbf{\theta}}}_{t}=h({\boldsymbol{\mathbf{\theta}}}_{t+1},{\boldsymbol{\mathbf{x}}}_{t-1},\alpha_{t}))\propto p_{\mathrm{I}}({\boldsymbol{\mathbf{x}}}_{t};{\boldsymbol{\mathbf{\theta}}}_{t+1})p_{\mathrm{S}}({\boldsymbol{\mathbf{x}}}_{t-1}\,|\,{\boldsymbol{\mathbf{x}}}_{t};\alpha_{t})$, where $h({\boldsymbol{\mathbf{\theta}}}_{t+1},{\boldsymbol{\mathbf{x}}}_{t-1},\alpha_{t})$ is the Bayesian update function. By feeding this posterior parameter ${\boldsymbol{\mathbf{\theta}}}_{t}$ into a neural network $\psi(\cdot)$, ${\boldsymbol{\mathbf{x}}}_{t}$’s output distribution $p_{\mathrm{O}}(\cdot)$ is parameterized as $p_{\mathrm{O}}({\boldsymbol{\mathbf{x}}}_{t};\psi({\boldsymbol{\mathbf{\theta}}}_{t}))$. Finally, a receiver distribution $p_{\mathrm{R}}(\cdot)$ is defined as the expectation of the sender distribution with respect to the output distribution, i.e., $p_{\mathrm{R}}({\boldsymbol{\mathbf{x}}}_{t-1};\psi({\boldsymbol{\mathbf{\theta}}}_{t}),\alpha_{t}):=\mathbb{E}_{p_{\mathrm{O}}({\boldsymbol{\mathbf{x}}}_{t};\psi({\boldsymbol{\mathbf{\theta}}}_{t}))}[p_{\mathrm{S}}({\boldsymbol{\mathbf{x}}}_{t-1}\,|\,{\boldsymbol{\mathbf{x}}}_{t};\alpha_{t})]$. See Figure 1 (a) for a visualization of the relationships between these distributions.
@@ 3 ParamReL: Parameter Space Representation Learning Framework
Prior for ${\boldsymbol{\mathbf{z}}}_{t}$: Ideally, the latent semantics $\{{\boldsymbol{\mathbf{z}}}_{t}\}_{t}$ should provide low-dimensional representations that are different from the intermediates $\{{\boldsymbol{\mathbf{x}}}_{t}\}_{t}$ of BFNs, yet they should not compromise the data reconstruction process. To generate high-quality semantic representations, a learnable and smooth latent space is essential [4]. This necessitates the integration of the prior distribution $p({\boldsymbol{\mathbf{z}}}_{t})$ into a robust probabilistic framework, thereby enabling the unconditional sampling of ${{\boldsymbol{\mathbf{x}}}}_{0}$. For simplicity and sampling efficiency, $p({\boldsymbol{\mathbf{z}}}_{t})$ could be assumed to follow a straightforward distribution, such as a Gaussian.
@@ 3.2 Training and Testing with ParamReL
Maximizing the objective function $\mathcal{L}_{\text{ParamReL}}$ is equivalent to performing amortized inference  [22] through encoders $q_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{z}}}_{t}|{\boldsymbol{\mathbf{\theta}}}_{t},t)$ and learning likelihood function through decoders [55]. The encodable posterior $q_{{\boldsymbol{\mathbf{\phi}}}}\left({\boldsymbol{\mathbf{z}}}_{t}|{\boldsymbol{\mathbf{\theta}}}_{t},t\right)$ is used to infer high-level semantics $\{{{\boldsymbol{\mathbf{x}}}}_{t}\}_{t}$. During the decoding stage, ${{\boldsymbol{\mathbf{x}}}}_{0}$ is used into generative latent variables ${\boldsymbol{\mathbf{x}}}_{1:T}$, which contains low-level semantic features in generating the observations. In ParamReL, the parameters of the output distribution are learned through iteratively proceeding the Bayesian updating process and a learned noise model ${\psi}({{\boldsymbol{\mathbf{\theta}}}},{\boldsymbol{\mathbf{z}}})$ parameterized by neural networks $\psi$.
@@ 3.3 Regularizing Semantics By Maximizing Mutual Information
Ideally, during the training phase, we want to acquire the semantic representation ${\boldsymbol{\mathbf{z}}}_{t}$ by the self-encoder $q_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{z}}}_{t}|{\boldsymbol{\mathbf{\theta}}}_{t},t)$ and achieve high-quality reconstruction $\widehat{{\boldsymbol{\mathbf{x}}}_{0}}$ by the decoder (i.e., the output distribution $p_{O}({\boldsymbol{\mathbf{x}}}_{0};\psi({\boldsymbol{\mathbf{\theta}}}_{0},{\boldsymbol{\mathbf{z}}}_{0}))$). However, there exists a trade-off between inference and learning [34, 46] coherent in optimizing the objective function $\mathcal{L}_{\text{ParamReL}}$ in Eq. (3). In most cases, ParamReL favours fitting likelihood rather than inference [56]. Based on the rate-distortion theory [1, 3], the rate (KL divergence term) constrained by the encoders compresses enough information for the distortion (reconstruction error) but less informative to induce a smooth latent space.
@@
Mutual Information Learning Unlike the rest of the terms that can be optimized directly using reparameterization tricks, the TC term cannot be directly optimized due to intractable marginal distribution $q_{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{z}}}_{t}})$. Here, we follow the guidance in  [56] to replace the TC term with any strict divergence $D$, where $D\left(q_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{z}}})\|p({\boldsymbol{\mathbf{z}}})\right)=0$ iff $q_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{z}}})=p({\boldsymbol{\mathbf{z}}})$. We implement the Maximum-Mean Discrepancy (MMD) [56] from the divergence family. MMD is a statistical measure that quantifies the difference between two probability distributions by comparing their mean embeddings in a high-dimensional feature space. By defining the kernel function $\kappa(\cdot,\cdot)$, $D_{\mathrm{MMD}}$ is denoted as:
@@ 4 Related Work
Recent advances have demonstrated that diffusion-based models [18, 37] are capable of generating high-quality data. Nonetheless, compared to the autoencoder framework, the intermediate outputs in diffusion stages are high-dimensional and lack smoothness, making them unsuitable for representation learning. Contemporary research focuses on encoding a conditional latent space to acquire semantic low-dimensional representations. However, sample-based models [31, 45], such as VAEs and Diffusions, exhibit limitations when applied to discrete data.
Deep hierarchical VAEs have seen progress in capturing latent dependence structures for encoding an expressive posterior, statistically or semantically. VQVAE-based  [44, 32] models have local-to-global feature-based explanatory hierarchies at the image level, forming a codebook-based discrete posterior. [36, 40] build recursive latent structures in multi-layer networks to form an aggregated posterior. NVAE [43] demonstrates that depth-wise hierarchies encoded by residual networks can approximate the posterior precisely despite using shallow networks. Unlike the sample-based encoder, where the information flow between input and latent is maximized in encoding-decoding pipelines on the sample space, we develop progressive encoders in the parameter space to capture the dynamic semantics.
@@ 5 Experiments
We present two ParamReL variants in different parameter spaces: for the continuous input distribution, denoted as ParamReLc, and for the discrete input distribution, denoted as ParamReLd; refer to Table  1 for the distribution definitions. We evaluate ParamReLs on discrete and continuous benchmarks, including three discrete datasets binarized by grey-style data: binarized MNIST (bMNIST) [8], binarized FashionMNIST (bFashionMNIST)  [47]; as well as three continuous datasets: CelebA  [27], CIFAR10  [24], and Shapes3d  [5]  [^2] For the discrete version, continuous data ($k$-bit images) can be discretized into $2^{k}$ bins. This process involves dividing the data range $[-1,1]$ into $k$ intervals, each with a length of $2/k$..The hyperparameter choices and experiment configurations corresponding to each dataset for training can be found in Appendix D.
@@
We conduct a comprehensive two-fold comparison. Firstly, we compare our parameter-based variants against sample-based representation learning baselines, including AE and VAE-based models: $\beta$-VAE [17], infoVAE [55], and diffusion-based models: DiffusionAE [31], infoDiffusion [45]. These models represent significant advancements in representation learning, with $\beta$-VAE being the pioneer in VAE-based disentanglement, infoVAE introducing MMD to strike a balance between generation and representation, DiffusionAE integrating AEs into Diffusion models for encodable latent learning, and infoDiffusion incorporating VAEs into diffusion models for disentanglement learning. Secondly, we compare the ParamReLd and ParamReLc variants across various input distributions for continuous and discrete data representation learning. This comparison enables a detailed investigation of parameter space assumptions for discrete and continuous representation learning.

### v3

@@ Abstract
The recently proposed Bayesian Flow Networks (BFNs) show great potential in modeling parameter spaces, offering a unified strategy for handling continuous, discretized, and discrete data. However, BFNs cannot learn high-level semantic representation from the parameter space since common encoders, which encode data into one static representation, cannot capture semantic changes in parameters. This motivates a new direction: learning semantic representations hidden in the parameter spaces to characterize mixed-typed noisy data. Accordingly, we propose a representation learning framework named ParamReL, which operates in the parameter space to obtain parameter-wise latent semantics that exhibit progressive structures. Specifically, ParamReL proposes a self-encoder to learn latent semantics directly from parameters, rather than from observations. The encoder is then integrated into BFNs, enabling representation learning with various formats of observations. Mutual information terms further promote the disentanglement of latent semantics and capture meaningful semantics simultaneously. We illustrate conditional generation and reconstruction in ParamReL via expanding BFNs, and extensive quantitative experimental results demonstrate the superior effectiveness of ParamReL in learning parameter representation. We release the code at: https://github.com/amasawa/ParamReL/tree/main.
@@ 1 Introduction
Representation learning bengio2013representation , which aims at learning low-dimensional latent semantics from high-dimensional observations, offers an unsupervised approach to discovering high-level semantics in observations. It has been widely applied in areas such as computer vision CVPRLiCMZKK23 ; CVPRZhaoWL23 ; CVPRDongHLLQG23 , and data analytics tonekaboni2022decoupling ; oublal2024disentangling . While most representation learning methods kim2018disentangling ; chen2018isolating ; meo2023alpha work on continuous-valued observations, different non-trivial methods are needed to discover semantics for the discretized  van2017neural ; razavi2019generating and discrete data austin2021structured ; chen2022analog . Consequently, these individual efforts might face issues such as inconsistent discoveries within the data zhou2024beta or repeated modelling efforts krishnan2018challenges ; zhao2020variational .
On the other hand, Bayesian Flow Networks (BFNs) graves2023bayesian ; song2024unified ; xue2024unifying have been recently proposed as a promising deep generative model. By operating in the parameter space, BFNs design a multi-step mechanism to approximate the ground-truth parameters of observation sequentially. As a result, a uniform strategy may be adopted to deal with continuous, discretized, and discrete data while simultaneously maintaining fast sampling. Pilot studies of BFNs have shown promising results in modelling different data formats.
@@ 2 Bayesian Flow Networks
Bayesian Flow Networks (BFNs) graves2023bayesian ; song2024unified ; xue2024unifying are a new class of latent variable models. Similar to the reverse process of diffusion models ho2020denoising ; song2020denoising , BFNs target optimizing the neural network architecture on posterior parameters to simulate observations as they are gradually observed. In particular, BFNs assume two types of distributions: a simple input distribution $P_{\mathrm{I}}(\cdot)$ representing the initial belief about observations and an output distribution $P_{\mathrm{O}}(\cdot)$ representing the simulating distribution for observations. The parameters of input distribution are first updated through a Bayesian inference scheme and then passed into a neural network to form the parameters of output distributions. The main objective of BFNs is to minimize the Kullback-Leibler (KL) divergence between the ground-truth data distribution and the output distribution, ensuring that the output distribution closely approximates the ground-truth data distribution.
@@
Following the notations in diffusion models, we use ${\boldsymbol{\mathbf{x}}}_{0}$ to denote the observations. There are $T$ reverse steps in BFNs which gradually reveals the information of ${\boldsymbol{\mathbf{x}}}_{0}$ through $\{{\boldsymbol{\mathbf{x}}}_{T},{\boldsymbol{\mathbf{x}}}_{T-1},\ldots,{\boldsymbol{\mathbf{x}}}_{1}\}$ to the input distribution [^1] It is noted that the index $t$ is used reversely in graves2023bayesian . We make such changes to be consistent with the diffusion model settings ho2020denoising ; song2020denoising .. At the $t$-th step, the reverse step is denoted as a sender distribution $p_{\mathrm{S}}({\boldsymbol{\mathbf{x}}}_{t-1}\,|\,{\boldsymbol{\mathbf{x}}}_{t};\alpha_{t})$, where $\alpha_{t}$ is a precision parameter. Combined with input distribution $p_{\mathrm{I}}({\boldsymbol{\mathbf{x}}}_{t};{\boldsymbol{\mathbf{\theta}}}_{t+1})$, the posterior distribution of ${\boldsymbol{\mathbf{x}}}_{t}$ is obtained as $p({\boldsymbol{\mathbf{x}}}_{t};{\boldsymbol{\mathbf{\theta}}}_{t}=h({\boldsymbol{\mathbf{\theta}}}_{t+1},{\boldsymbol{\mathbf{x}}}_{t-1},\alpha_{t}))\propto p_{\mathrm{I}}({\boldsymbol{\mathbf{x}}}_{t};{\boldsymbol{\mathbf{\theta}}}_{t+1})p_{\mathrm{S}}({\boldsymbol{\mathbf{x}}}_{t-1}\,|\,{\boldsymbol{\mathbf{x}}}_{t};\alpha_{t})$, where $h({\boldsymbol{\mathbf{\theta}}}_{t+1},{\boldsymbol{\mathbf{x}}}_{t-1},\alpha_{t})$ is the Bayesian update function. By feeding this posterior parameter ${\boldsymbol{\mathbf{\theta}}}_{t}$ into a neural network $\psi(\cdot)$, ${\boldsymbol{\mathbf{x}}}_{t}$’s output distribution $p_{\mathrm{O}}(\cdot)$ is parameterized as $p_{\mathrm{O}}({\boldsymbol{\mathbf{x}}}_{t};\psi({\boldsymbol{\mathbf{\theta}}}_{t}))$. Finally, a receiver distribution $p_{\mathrm{R}}(\cdot)$ is defined as the expectation of the sender distribution with respect to the output distribution, i.e., $p_{\mathrm{R}}({\boldsymbol{\mathbf{x}}}_{t-1};\psi({\boldsymbol{\mathbf{\theta}}}_{t}),\alpha_{t}):=\mathbb{E}_{p_{\mathrm{O}}({\boldsymbol{\mathbf{x}}}_{t};\psi({\boldsymbol{\mathbf{\theta}}}_{t}))}[p_{\mathrm{S}}({\boldsymbol{\mathbf{x}}}_{t-1}\,|\,{\boldsymbol{\mathbf{x}}}_{t};\alpha_{t})]$. See Figure 1 (a) for a visualization of the relationships between these distributions.
@@ 3 ParamReL: Parameter Space Representation Learning Framework
Prior for ${\boldsymbol{\mathbf{z}}}_{t}$: Ideally, the latent semantics $\{{\boldsymbol{\mathbf{z}}}_{t}\}_{t}$ should provide low-dimensional representations that are different from the intermediates $\{{\boldsymbol{\mathbf{x}}}_{t}\}_{t}$ of BFNs, yet they should not compromise the data reconstruction process. To generate high-quality semantic representations, a learnable and smooth latent space is essential bengio2013representation . This necessitates the integration of the prior distribution $p({\boldsymbol{\mathbf{z}}}_{t})$ into a robust probabilistic framework, thereby enabling the unconditional sampling of ${{\boldsymbol{\mathbf{x}}}}_{0}$. For simplicity and sampling efficiency, $p({\boldsymbol{\mathbf{z}}}_{t})$ could be assumed to follow a straightforward distribution, such as a Gaussian.
@@ 3.2 Training and Testing with ParamReL
Maximizing the objective function $\mathcal{L}_{\text{ParamReL}}$ is equivalent to performing amortized inference  kingma2013auto through encoders $q_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{z}}}_{t}|{\boldsymbol{\mathbf{\theta}}}_{t},t)$ and learning likelihood function through decoders zhao2017infovae . The encodable posterior $q_{{\boldsymbol{\mathbf{\phi}}}}\left({\boldsymbol{\mathbf{z}}}_{t}|{\boldsymbol{\mathbf{\theta}}}_{t},t\right)$ is used to infer high-level semantics $\{{{\boldsymbol{\mathbf{x}}}}_{t}\}_{t}$. During the decoding stage, ${{\boldsymbol{\mathbf{x}}}}_{0}$ is used into generative latent variables ${\boldsymbol{\mathbf{x}}}_{1:T}$, which contains low-level semantic features in generating the observations. In ParamReL, the parameters of the output distribution are learned through iteratively proceeding the Bayesian updating process and a learned noise model ${\psi}({{\boldsymbol{\mathbf{\theta}}}},{\boldsymbol{\mathbf{z}}})$ parameterized by neural networks $\psi$.
@@ 3.3 Regularizing Semantics By Maximizing Mutual Information
Ideally, during the training phase, we want to acquire the semantic representation ${\boldsymbol{\mathbf{z}}}_{t}$ by the self-encoder $q_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{z}}}_{t}|{\boldsymbol{\mathbf{\theta}}}_{t},t)$ and achieve high-quality reconstruction $\widehat{{\boldsymbol{\mathbf{x}}}_{0}}$ by the decoder (i.e., the output distribution $p_{O}({\boldsymbol{\mathbf{x}}}_{0};\psi({\boldsymbol{\mathbf{\theta}}}_{0},{\boldsymbol{\mathbf{z}}}_{0}))$). However, there exists a trade-off between inference and learning shao2020controlvae ; wu2023evae coherent in optimizing the objective function $\mathcal{L}_{\text{ParamReL}}$ in Eq. (3). In most cases, ParamReL favours fitting likelihood rather than inference zhao2019infovae . Based on the rate-distortion theory alemi2018fixing ; bae2022multi , the rate (KL divergence term) constrained by the encoders compresses enough information for the distortion (reconstruction error) but less informative to induce a smooth latent space.
@@
Mutual Information Learning Unlike the rest of the terms that can be optimized directly using reparameterization tricks, the TC term cannot be directly optimized due to intractable marginal distribution $q_{{\boldsymbol{\mathbf{\phi}}}}({{\boldsymbol{\mathbf{z}}}_{t}})$. Here, we follow the guidance in  zhao2019infovae to replace the TC term with any strict divergence $D$, where $D\left(q_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{z}}})\|p({\boldsymbol{\mathbf{z}}})\right)=0$ iff $q_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{z}}})=p({\boldsymbol{\mathbf{z}}})$. We implement the Maximum-Mean Discrepancy (MMD) zhao2019infovae from the divergence family. MMD is a statistical measure that quantifies the difference between two probability distributions by comparing their mean embeddings in a high-dimensional feature space. By defining the kernel function $\kappa(\cdot,\cdot)$, $D_{\mathrm{MMD}}$ is denoted as:
@@ 4 Related Work
Recent advances have demonstrated that diffusion-based models ho2020denoising ; song2020denoising are capable of generating high-quality data. Nonetheless, compared to the autoencoder framework, the intermediate outputs in diffusion stages are high-dimensional and lack smoothness, making them unsuitable for representation learning. Contemporary research focuses on encoding a conditional latent space to acquire semantic low-dimensional representations. However, sample-based models preechakul2022diffusion ; wang2023infodiffusion , such as VAEs and Diffusions, exhibit limitations when applied to discrete data.
Deep hierarchical VAEs have seen progress in capturing latent dependence structures for encoding an expressive posterior, statistically or semantically. VQVAE-based  van2017neural ; razavi2019generating models have local-to-global feature-based explanatory hierarchies at the image level, forming a codebook-based discrete posterior. sonderby2016ladder ; tomczak2018vae build recursive latent structures in multi-layer networks to form an aggregated posterior. NVAE vahdat2020nvae demonstrates that depth-wise hierarchies encoded by residual networks can approximate the posterior precisely despite using shallow networks. Unlike the sample-based encoder, where the information flow between input and latent is maximized in encoding-decoding pipelines on the sample space, we develop progressive encoders in the parameter space to capture the dynamic semantics.
@@ 5 Experiments
We present two ParamReL variants in different parameter spaces: for the continuous input distribution, denoted as ParamReLc, and for the discrete input distribution, denoted as ParamReLd; refer to Table  1 for the distribution definitions. We evaluate ParamReLs on discrete and continuous benchmarks, including three discrete datasets binarized by grey-style data: binarized MNIST (bMNIST) deng2012mnist , binarized FashionMNIST (bFashionMNIST)  xiao2017fashion ; as well as three continuous datasets: CelebA  liu2015deep , CIFAR10  krizhevsky2009learning , and Shapes3d  3dshapes18  [^2] For the discrete version, continuous data ($k$-bit images) can be discretized into $2^{k}$ bins. This process involves dividing the data range $[-1,1]$ into $k$ intervals, each with a length of $2/k$..The hyperparameter choices and experiment configurations corresponding to each dataset for training can be found in Appendix D.
@@
We conduct a comprehensive two-fold comparison. Firstly, we compare our parameter-based variants against sample-based representation learning baselines, including AE and VAE-based models: $\beta$-VAE higgins2017beta , infoVAE zhao2017infovae , and diffusion-based models: DiffusionAE preechakul2022diffusion , infoDiffusion wang2023infodiffusion . These models represent significant advancements in representation learning, with $\beta$-VAE being the pioneer in VAE-based disentanglement, infoVAE introducing MMD to strike a balance between generation and representation, DiffusionAE integrating AEs into Diffusion models for encodable latent learning, and infoDiffusion incorporating VAEs into diffusion models for disentanglement learning. Secondly, we compare the ParamReLd and ParamReLc variants across various input distributions for continuous and discrete data representation learning. This comparison enables a detailed investigation of parameter space assumptions for discrete and continuous representation learning.